OpenStudy (anonymous):

*write and solve an algebraic equation* two angles form a linear pair, the measure of one angle is 15(degrees) less than 2/3 the measure of one angle

OpenStudy (tkhunny):

You should start! A = Measure of First Angle B = Measure of Second Angle Now that we've done that very important step - the definition, let's write equations from sentences... "two angles form a linear pair" Okay, how do you write that in terms of A and B? "the measure of one angle is 15(degrees) less than 2/3 the measure of [the other] angle" Super. How do you write that in terms of A and B? Go!!

OpenStudy (anonymous):

makes no sense but k

OpenStudy (tkhunny):

What makes no sense? The Definitions? That's the most important part.

OpenStudy (anonymous):

just gonna search it up and write down random answer pce@!!!!!!!!

OpenStudy (tkhunny):

Probably not going to work for you on an exam. Good luck.

OpenStudy (anonymous):

117 and 63 degrees. \[\left(\frac{2 \theta }{3}-15\right)+\theta =180 \]

OpenStudy (tkhunny):

Or, as robtobey has demonstrated, there is no need to be organized and exact and clear if you can just pull it out of the air. I'm guessing that won't help you much on an exam, either.